# Filtrations on derived category of twisted K3 surfaces

### v1

@@ 4.1. Construction of $\mathfrak{o}_{\mathscr{X}}$
When $\mathscr{X}\to X$ is trivial, this is exactly [12, Lemma 2.14]. In the case $\mathscr{X}\to X$ is essentially trivial, after taking a twisted line bundle $\mathcal{L}\in\operatorname{Pic}(\mathscr{X})$, we can identify $\operatorname{Hom}_{\mathcal{O}_{\mathscr{X}}}(\mathcal{E},\mathcal{F})$ with $\operatorname{Hom}_{\mathcal{O}_{X}}(\mathcal{E}\otimes\mathcal{L}^{\vee},\mathcal{F}\otimes\mathcal{L}^{\vee})$ and their schematic degenerate loci are the same. The assertion follows from the untwisted version.
@@
For simplicity, we assume that $\mathscr{M}_{i}=M_{i}$ and $\mathscr{N}_{j}=N_{j}$ are fine moduli spaces. Otherwise, one may either work with everything over stacks directly or pass to the associated Severi-Brauer varieties just as in the proof of [15, Theorem 3.1]. Take an irreducible component $Y_{i}\subseteq O_{\mathcal{E}_{i}}$ of dimension $n_{i}$ and $Z_{j}\subseteq O_{\mathcal{F}_{j}}$ of dimension $m_{j}$, we set
@@
[…]
@@
Due to [19, Theorem], $\mathrm{Image}(\psi)$ is a constant cycle subvariety. By Theorem 3.8, we get
@@
Let $\mathcal{E}\in\mathrm{D}^{b}(\mathscr{X})$. As the argument proceeds a similar way as in [25, 21], we only sketch the proof.
@@
    The existence of such vectors is ensured by Remark 4.4. By Proposition 4.3, there exist $\mathscr{X}$-twisted locally free sheaves
@@
where $\mathcal{G}$ is the shifted sheaf $h^{m}(\mathcal{E})[-m]$ and $\mathcal{F}\in\mathrm{D}^{b}(X)$ is the truncated complex $\tau^{\leq m-1}$ By induction on the length of $\mathcal{E}$, one can get $c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X)$. ∎
4.12. SYZ filtration v.s. Voisin’s filtration
Now we can generalize the work in [21] and [15] to twisted K3 surfaces. With the twisted Beauville-Voisin class $\mathfrak{o}_{\mathscr{X}}$, we are allowed to define the twisted incidence correspondence
$\Gamma^{\prime}:=\Big{\{}(\xi,\mathcal{E})~{}|~{}c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}=[\xi]+k\mathfrak{o}_{X}\in\operatorname{CH}_{0}(X)\Big{\}}\subset X^{[d]}\times M_{\sigma}(\mathscr{X},\mathbf{v}),$ (33)
for arbitrary $\mathbf{v}$. As before, due to Theorem 1.2, the projection maps $\pi_{1}:\Gamma^{\prime}\to M_{\sigma}(\mathscr{X},\mathbf{v})$ and $\pi_{2}:\Gamma^{\prime}\to X^{[n]}$ are dominant.
So we can obtain a subvariety $\Gamma\subseteq\Gamma^{\prime}$ such that $\Gamma$ is generically finite over $M_{\sigma}(\mathscr{X},\mathbf{v})$ and $X^{[d]}$. For two general points $q,p$ in $\Gamma$ lying on the same fiber of $\pi_{1}$ (resp. $\pi_{2}$), we have $\pi_{2}(p)=\pi_{2}(q)\in\operatorname{CH}_{0}(X^{[d]})$ (resp. $\pi_{1}(p)=\pi_{1}(q)\in\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))$) by Theorem 2.6. This implies that if $Z\subset X^{[d]}$ is a constant cycle subvariety, $\pi_{1}(\pi_{2}^{-1}(Z))$ is a constant cycle subvariety as well.
Theorem 4.13.
If $c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{Y}}\in\mathbf{S}_{i}(X)$, then $\dim O_{\mathcal{E}}\geq d-i.$ In other words, there is an inclusion
$\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))\subseteq\mathbf{S}_{\bullet}^{\operatorname{BV}}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v})).$
Conversely, if $\dim O_{\mathcal{E}}\cap M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})\geq d-i$, then $\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X}))$. In particular, if $M_{\sigma}(\mathscr{X},\mathbf{v})$ consists of locally free sheaves, then
$\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))=\mathbf{S}_{\bullet}^{\operatorname{BV}}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v})).$
Proof.
The proof is identical to the proof of [21, Theorem 0.5]. ∎
Corollary 4.14.
For any $\mathcal{E}\in M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})$ with dim $\dim O_{\mathcal{E}}\cap M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})=d-i$, the subvarieties of the orbit $O_{\mathcal{E}}\cap M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})$ with dimension $d-i$ are dense in $M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})$.
Proof.
This can be obtained along the same lines in the proof [15, Corollary 3.2]. ∎
4.15. Further remarks
If there is a $\mathcal{F}_{0}$ with $c_{2}(\mathcal{F}_{0})\in\mathbf{S}_{0}(A)$, one can thus define a Shen-Yin-Zhao type filtration on $K_{\sigma}(\mathscr{A},\mathbf{v})$. In this situation, our construction of (twisted) incidence correspondences is also valid and one can obtain a similar result for the degenerate loci of stable locally free sheaves on twisted abelian surfaces. The main problem will be the very first step, that is to check whether the induced SYZ-type filtration coincide with Voisin’s filtration when $K=\ker(A^{[n+1]}\to A)$. Note that in this case, one does have an element $\mathcal{F}_{0}\in K$ with $c_{2}(\mathcal{F}_{0})\in\mathbf{S}_{0}(A)$.
5. Proof of Theorem 1.4
@@ 5.1. Beauville-Voisin subgroup
Let $\dim M_{\sigma}(\mathscr{X},v)=2n$. Note that $M_{\sigma}(\mathscr{X},\mathbf{v})$ admits a birational Lagrangian fibration. As proved in Proposition 3.2, there exists $\mathcal{E}\in M_{\sigma(\mathscr{X},\mathbf{v})}$ such that $O_{\mathcal{E}}$ has dimension $n$ and its maximal dimensional irreducible components are Zariski dense in $M_{\sigma}(\mathscr{X},\mathbf{v})$. Set $\mathcal{F}=\Psi(\mathcal{E})\in\mathrm{D}^{b}(\mathscr{Y})$ and $\mathbf{w}=\mathbf{v}(\mathcal{F})$. The derived equivalence $\Psi$ induces an isomorphism
@@
In [21, Question 4.1], it has been asked if the map
@@
$\mathcal{E}^{\prime}\cong\mathbb{C}_{x_{1}}\oplus\cdots\oplus\mathbb{C}_{x_{i}}\in\mathrm{D}^{b}(\mathscr{X}).$
@@
By Theorem 1.2, $\Psi(\mathcal{E}^{\prime})\in\mathbf{S}_{\mathrm{d}(\mathcal{E}^{\prime})}(\mathrm{D}^{b}(\mathscr{Y}))=\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{Y}))$. Hence $\Psi(\mathcal{E})\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{Y}))$ as well. ∎
@@ 6.1. Bloch’s conjecture for zero cycles
[…]
@@
Conjecture 6.2 (Bloch’s conjecture for Bridgeland moduli space).
Let $Y$ be a Bridgeland moduli space of dimension $2n$. Set
$\operatorname{Gr}_{s}\operatorname{CH}_{0}(Y):=\mathbf{S}_{s}^{\rm SYZ}\operatorname{CH}_{0}(Y)/\mathbf{S}_{s-1}^{\rm SYZ}\operatorname{CH}_{0}(Y)$
@@
This conjecture has been studied in [14] for untwisted Bridgeland moduli spaces, i.e. when $Y$ is isomorphic to the moduli space of stable objects in $\mathrm{D}^{b}(X)$ for some K3 surface $X$. The method in [14] is also valid for twisted Bridgeland moduli spaces once there is a “right” formulation of $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(Y)$. We will explain this in the next subsection.
@@ 6.3. Birational motives
We denote by $\mathfrak{h}^{\circ}(Y)$ the birational motive of $Y$ in the sense of [13]. It has been conjectured in [22] that there is a natural grading on $\mathfrak{h}^{\circ}(Y)$ compatible with the co-algebra structure on $\mathfrak{h}^{0}(Y)$ (cf[22, Conjecture 2]). When $Y=M_{\sigma}(X,\mathbf{v})$ for untwisted K3, such grading is governed by the isomorphism
@@
as co-algebra objects by [22, Theorem 4], where $\dim Y=2n$. The grading on $\mathfrak{h}^{\circ}(Y)$ is compatible with $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(Y)$, i.e.
@@
(cf[22, Section 6.4 and Theorem 7.3]). Due to Theorem 1.2, we can provide new examples for [22, Conjecture 2]. The idea is that the Shen–Yin–Zhao type filtration admits a “motivic” splitting on $\mathfrak{h}^{\circ}(Y)$.
@@
As shown before, $\Gamma\subseteq Y\times X^{[n]}$ is a generically finite and dominant correspondence. The assertion is then a direct consequence of [22, Proposition 2.3]. ∎
@@
By Proposition 6.4, we only need to treat the case $M_{\sigma}(\mathscr{X},\mathbf{v})\cong X^{[n]}$. Then the result follows from [14, Proposition 3.8(iii)]. ∎
@@
By [1, Proposition 4][10, Lemma 2.6](see also [14, Corollary 3.3]), $Y\cong M_{\sigma}(\mathscr{X},\mathbf{v})$ for some twisted K3 surface $\mathscr{X}\to X$ and hence carries a Shen-Yin-Zhao type filtration on $\operatorname{CH}_{0}(Y)$.

### v3

@@ 4.1. Construction of $\mathfrak{o}_{\mathscr{X}}$
When $\mathscr{X}\to X$ is trivial, this is exactly [14, Lemma 2.14]. In the case $\mathscr{X}\to X$ is essentially trivial, after taking a twisted line bundle $\mathcal{L}\in\operatorname{Pic}(\mathscr{X})$, we can identify $\operatorname{Hom}_{\mathcal{O}_{\mathscr{X}}}(\mathcal{E},\mathcal{F})$ with $\operatorname{Hom}_{\mathcal{O}_{X}}(\mathcal{E}\otimes\mathcal{L}^{\vee},\mathcal{F}\otimes\mathcal{L}^{\vee})$ and their schematic degeneracy loci are the same. The assertion follows from the untwisted version.
@@
For simplicity, we assume that $\mathscr{M}_{i}=M_{i}$ and $\mathscr{N}_{j}=N_{j}$ are fine moduli spaces. Otherwise, one may either work with everything over stacks directly or pass to the associated Severi-Brauer varieties just as in the proof of [17, Theorem 3.1]. Take an irreducible component $Y_{i}\subseteq O_{\mathcal{E}_{i}}$ of dimension $n_{i}$ and $Z_{j}\subseteq O_{\mathcal{F}_{j}}$ of dimension $m_{j}$, we set
@@
${\mathbf{P}}$${M_{H}(\mathscr{X},\mathbf{v}_{0})}$${Y\times Z}$$\scriptstyle{\psi}$$\scriptstyle{\pi}$
@@
Due to Theorem 2.3, $\mathrm{Image}(\psi)$ is a constant cycle subvariety. By Theorem 3.5, we get
@@
Let $\mathcal{E}\in\mathrm{D}^{(1)}(\mathscr{X})$. As the argument proceeds a similar way as in [27, 23], we only sketch the proof.
@@
    By Proposition 4.3, there exist $\mathscr{X}$-twisted locally free sheaves
@@
where $\mathcal{G}$ is the shifted sheaf $h^{m}(\mathcal{E})[-m]$ and $\mathcal{F}\in\mathrm{D}^{(1)}(\mathscr{X})$ is the truncated complex $\tau^{\leq m-1}$ By induction on the length of $\mathcal{E}$, one can get $c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X)$. ∎
5. Proof of Main Theorems
@@ 5.1. Beauville-Voisin subgroup
Let $\dim M_{\sigma}(\mathscr{X},v)=2n$. Note that $M_{\sigma}(\mathscr{X},\mathbf{v})$ admits a birational Lagrangian fibration. As proved in Proposition 3.1, there exists $\mathcal{E}\in M_{\sigma(\mathscr{X},\mathbf{v})}$ such that $O_{\mathcal{E}}$ has dimension $n$ and its maximal dimensional irreducible components are Zariski dense in $M_{\sigma}(\mathscr{X},\mathbf{v})$. Set $\mathcal{F}=\Psi(\mathcal{E})\in\mathrm{D}^{(1)}(\mathscr{Y})$ and $\mathbf{w}=\mathbf{v}(\mathcal{F})$. The derived equivalence $\Psi$ induces an isomorphism
@@
In [23, Question 4.1], it has been asked if the map
@@
$\mathcal{E}^{\prime}\cong\mathbb{C}_{x_{1}}\oplus\cdots\oplus\mathbb{C}_{x_{i}}\in\mathrm{D}^{(1)}(\mathscr{X}).$
@@
By Theorem 1.1, $\Psi(\mathcal{E}^{\prime})\in\mathbf{S}_{\mathrm{d}(\mathcal{E}^{\prime})}(\mathrm{D}^{(1)}(\mathscr{Y}))=\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{Y}))$. Hence $\Psi(\mathcal{E})\in\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{Y}))$ as well. ∎
5.3. SYZ filtration v.s. Voisin’s filtration
We now generalize the work in [23] and [17] to twisted K3 surfaces. Assume that $M_{\sigma}(\mathscr{X},\mathbf{v})$ is a moduli space of $\sigma$-stable locally free sheaves on $\mathscr{X}$ and $\sigma$ is $\mathbf{v}$-generic. Using the twisted Beauville-Voisin class $\mathfrak{o}_{\mathscr{X}}$, we define the twisted incidence correspondence:
$\Gamma^{\prime}:=\left\{(\xi,\mathcal{E}):c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}=[\xi]+k\mathfrak{o}_{X}\in\operatorname{CH}_{0}(X)\right\}\subset X^{[d]}\times M_{\sigma}(\mathscr{X},\mathbf{v}),$ (37)
for an arbitrary Mukai vector $\mathbf{v}$. By Theorem 1.1, the projection maps $\pi_{1}\colon\Gamma^{\prime}\to M_{\sigma}(\mathscr{X},\mathbf{v})$ and $\pi_{2}\colon\Gamma^{\prime}\to X^{[d]}$ are both dominant.
Thus, there exists a subvariety $\Gamma\subseteq\Gamma^{\prime}$ that is generically finite over both $M_{\sigma}(\mathscr{X},\mathbf{v})$ and $X^{[d]}$. For two general points $p,q\in\Gamma$ lying on the same fiber of $\pi_{1}$ (respectively $\pi_{2}$), Theorem 2.3 implies that $\pi_{2}(p)=\pi_{2}(q)\in\operatorname{CH}_{0}(X^{[d]})$ (respectively $\pi_{1}(p)=\pi_{1}(q)\in\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))$). Consequently, if $Z\subset X^{[d]}$ is a constant cycle subvariety, then $\pi_{1}(\pi_{2}^{-1}(Z))$ is also constant cycle. Using the construction above, we obtain
Theorem 5.3.
If $c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{Y}}\in\mathbf{S}_{i}(X)$, then $\dim O_{\mathcal{E}}\geq d-i.$ Conversely, if $\dim O_{\mathcal{E}}\geq d-i$, then $\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{(1)}(\mathscr{X}))$. Consequently, we have
$\mathbf{S}^{\mathrm{SYZ}}_{\bullet}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))=\mathbf{S}^{\mathrm{BV}}_{\bullet}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v})).$
Proof.
By Theorems 1.3 and 7.2, we may assume (after applying a derived equivalence) that $M_{\sigma}(\mathscr{X},\mathbf{v})$ consists of locally free twisted sheaves. The remainder of the proof is identical to [23, Theorem 0.5]. ∎
Corollary 5.4.
For any $\mathcal{E}\in M_{\sigma}(\mathscr{X},\mathbf{v})$ with dim $\dim O_{\mathcal{E}}=d-i$, the subvarieties of the orbit $O_{\mathcal{E}}$ with dimension $d-i$ are dense in $M_{\sigma}(\mathscr{X},\mathbf{v})$.
Proof.
Again, we may assume $M_{\sigma}(\mathscr{X},\mathbf{v})$ consists of locally free twisted sheaves. Then this can be obtained along the same lines as in the proof [17, Corollary 3.2]. ∎
5.4. Proof of Corollary 1.5
By [1, Proposition 4], [10, Lemma 2.6], and [16, Corollary 3.3], the condition on the Markman-Mukai lattice implies that $Y\cong M_{\sigma}(\mathscr{X},\mathbf{v})$ for some twisted K3 surface $\mathscr{X}\to X$. Theorem 1.4 establishes:
$\mathbf{S}^{\mathrm{BV}}_{\bullet}\operatorname{CH}_{0}(Y)=\mathbf{S}^{\mathrm{SYZ}}_{\bullet}\operatorname{CH}_{0}(Y).$
To prove birational invariance, we adapt the argument in [17, Corollary 3.3] to the twisted case. Since the original proof omits some technical details, we provide a comprehensive exposition below.
By [2, Theorem 1.2], the hyper-Kähler birational models of $Y$ correspond to chambers in $\operatorname{Stab}^{\dagger}(\mathscr{X})$. As shown in [17, Corollary 3.3], any birational model $Y^{\prime}$ is isomorphic to $M_{\sigma^{\prime}}(\mathscr{X},\mathbf{v})$ for some generic $\sigma^{\prime}$ in a chamber $\mathcal{C}^{\prime}\subseteq\operatorname{Stab}^{\dagger}(\mathscr{X})$, and the birational map
$f:M_{\sigma}(\mathscr{X},\mathbf{v})\dashrightarrow M_{\sigma^{\prime}}(\mathscr{X},\mathbf{v})$
is induced by a derived (anti-)autoequivalence $\Phi\in\operatorname{Aut}(\mathrm{D}^{(1)}(\mathscr{X}))$ (up to automorphism).
Let $U\subset Y$ be the domain where $f$ is defined. For any $E\in U$, Theorem 1.3 guarantees:
$E\in\mathbf{S}_{i}^{\rm SYZ}(\mathrm{D}^{(1)}(\mathscr{X}))\implies\Phi(E)\in\mathbf{S}_{i}^{\rm SYZ}(\mathrm{D}^{(1)}(\mathscr{X}))$
F since derived equivalences preserve the SYZ filtration. Thus $f(E)\in\mathbf{S}_{i}^{\rm SYZ}\operatorname{CH}_{0}(Y^{\prime})$. Now consider any point $F$ with $[F]\in\mathbf{S}_{i}^{\operatorname{SYZ}}\operatorname{CH}_{0}(Y)$. By definition of the filtration, we have
$[F]=\sum n_{j}[F_{j}]\quad\text{with}\quad c_{2}(F_{j})\in\mathbf{S}_{i}(\mathscr{X}).$ (38)
By Theorem 2.3, the class $[F_{j}]$ is determined by $c_{2}(F_{j})$. Using the surjective projections $\pi_{1}:\Gamma^{\prime}\to M_{\sigma}(\mathscr{X},\mathbf{v})$ and $\pi_{2}:\Gamma^{\prime}\to X^{[d]}$ from the twisted incidence correspondence (37), we may adjust the $F_{j}$ to lie in $U$ while preserving (38).
Applying $f_{*}$ we obtain:
$f_{*}[F]=\sum n_{j}[f(F_{j})]\quad\text{with}\quad f(F_{j})\in\mathbf{S}_{i}^{\operatorname{SYZ}}\operatorname{CH}_{0}(Y^{\prime}).$
Hence $f_{*}[F]\in\mathbf{S}_{i}^{\operatorname{SYZ}}\operatorname{CH}_{0}(Y^{\prime})$. The reverse inclusion follows by applying the same argument to $f^{-1}$. We conclude:
$f_{*}(\mathbf{S}_{\bullet}^{\operatorname{SYZ}}\operatorname{CH}_{0}(Y))=\mathbf{S}_{\bullet}^{\operatorname{SYZ}}\operatorname{CH}_{0}(Y^{\prime}).$
Theorem 1.4 establishes $\mathbf{S}_{\bullet}^{\mathrm{BV}}=\mathbf{S}_{\bullet}^{\mathrm{SYZ}}$, completing the proof of birational invariance.
For the last assertion, note that if $\operatorname{rank}\operatorname{Pic}(Y)\notin\{2,3\}$, then either the algebraic Markman-Mukai lattice of $Y$ (which has rank $\geq 5$) contains an isotropic vector, or $Y$ admits no nontrivial birational hyper-Kähler model. The assertion follows immediately. ∎
5.5. Further remarks
If there is a $\mathcal{F}_{0}$ with $c_{2}(\mathcal{F}_{0})\in\mathbf{S}_{0}(A)$, one can thus define a Shen-Yin-Zhao type filtration on $K_{\sigma}(\mathscr{A},\mathbf{v})$. In this situation, our construction of (twisted) incidence correspondences is also valid and one can obtain a similar result for the degeneracy loci of stable locally free sheaves on twisted abelian surfaces. The main problem will be the very first step, that is to check whether the induced SYZ-type filtration coincides with Voisin’s filtration when $K=\ker(A^{[n+1]}\to A)$. Note that in this case, one does have an element $\mathcal{F}_{0}\in K$ with $c_{2}(\mathcal{F}_{0})\in\mathbf{S}_{0}(A)$.
@@ 6.1. Bloch’s conjecture for zero cycles
$\mathbf{S}_{\bullet}\operatorname{CH}_{0}(Y)=\mathbf{S}_{\bullet}^{\rm BV}\operatorname{CH}_{0}(Y).$ (41)
@@
Conjecture 6.1 (Bloch’s conjecture for (anti)-symplectic birational map).
Let $Y$ be a hyper-Kähler variety of dimension $2n$. Set
$\operatorname{Gr}_{s}\operatorname{CH}_{0}(Y):=\mathbf{S}_{s}^{\rm BV}\operatorname{CH}_{0}(Y)/\mathbf{S}_{s-1}^{\rm BV}\operatorname{CH}_{0}(Y)$
@@
This conjecture has been studied in [16] for untwisted Bridgeland moduli spaces, i.e. when $Y$ is isomorphic to the moduli space of stable objects in $\mathrm{D}^{b}(X)$ for some K3 surface $X$. The method in [16] is also valid for twisted Bridgeland moduli spaces. We will explain this in the next subsection.
@@ 6.2. Birational motives
We denote by $\mathfrak{h}^{\circ}(Y)$ the birational motive of $Y$ in the sense of [15]. It has been conjectured in [24] that there is a natural grading on $\mathfrak{h}^{\circ}(Y)$ compatible with the co-algebra structure on $\mathfrak{h}^{0}(Y)$ (cf[24, Conjecture 2]). When $Y=M_{\sigma}(X,\mathbf{v})$ for untwisted K3, such grading is governed by the isomorphism
@@
as co-algebra objects by [24, Theorem 4], where $\dim Y=2n$. The grading on $\mathfrak{h}^{\circ}(Y)$ is compatible with $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(Y)$, i.e.
@@
(cf[24, Section 6.4 and Theorem 7.3]). Due to Theorem 1.1, we can provide new examples for [24, Conjecture 2]. The idea is that the Shen–Yin–Zhao type filtration admits a “motivic” splitting on $\mathfrak{h}^{\circ}(Y)$.
@@
As shown before, $\Gamma\subseteq Y\times X^{[n]}$ is a generically finite and dominant correspondence. The assertion is then a direct consequence of [24, Proposition 2.3]. ∎
@@
By Proposition 6.2, we only need to treat the case $M_{\sigma}(\mathscr{X},\mathbf{v})\cong X^{[n]}$. Then the result follows from [16, Proposition 3.8(iii)]. ∎
@@
As discussed in the proof Corollary 1.5, note that $Y$ admits a birational Lagrangian fibration, this implies:
$Y\cong M_{\sigma}(\mathscr{X},\mathbf{v})$
for some twisted K3 surface $\mathscr{X}\to X$, which endows $\operatorname{CH}_{0}(Y)$ with the Shen-Yin-Zhao filtration $\mathbf{S}^{\mathrm{SYZ}}_{\bullet}\operatorname{CH}_{0}(Y)=\mathbf{S}^{\mathrm{BV}}_{\bullet}\operatorname{CH}_{0}(Y)$. We prove both implications for $\phi\in\mathrm{Bir}(Y)$ preserving the birational Lagrangian fibration:
  1. ($\Rightarrow$)
    When $\phi$ is (anti-)symplectic ($\phi^{*}|_{\mathrm{H}^{2,0}(Y)}=\pm\operatorname{id}$), [16, Theorem 2.17] implies $\phi_{*}$ acts as ±$\pm\operatorname{id}$ on $\mathrm{Gr}_{1}\operatorname{CH}_{0}(Y)$. Lemma 6.3 then yields the required action on higher graded pieces:
    $\phi_{*}|_{\mathrm{Gr}_{s}\operatorname{CH}_{0}(Y)}=(\pm 1)^{s}\operatorname{id}\quad\forall 1\leq s\leq n$
  2. ($\Leftarrow$)